OpenStudy (anonymous):

PLEASE HELLP????? the function f(x)=x^2. the graph of g(x) is f(x) translated to the right 3 unites and down 3 units. what is the function rule for g(x)?

OpenStudy (anonymous):

@nubeer @agent0smith can anyone help :(

OpenStudy (anonymous):

@jim_thompson5910 can you help??

terenzreignz (terenzreignz):

Okay, let's put it like this, first start with f(x): \[\Large y=x^2\]

OpenStudy (anonymous):

okay wait what..

terenzreignz (terenzreignz):

Sorry.... I don't know what's happening....

terenzreignz (terenzreignz):

\[\Large y = x^2\]

terenzreignz (terenzreignz):

crap, LaTeX isn't working hang on...

OpenStudy (anonymous):

okay haha.

terenzreignz (terenzreignz):

Okay, might as well make do with what I have -_- Anyway... start with y = x^2 yeah?

terenzreignz (terenzreignz):

If you want to translate up (or down) you manipulate the y-part of the equation. Say, if I want to translate it 7 units up, it would be y - 7 = x^2 Get it?

OpenStudy (anonymous):

okay yeah kinda haha

OpenStudy (anonymous):

so since im going 3 down itd be y-3=x^2??

terenzreignz (terenzreignz):

No.... - (strangely enough) is for going UP + is for going down. Don't ask me why, it's going to take a nauseating geometric explanation... (although go ahead and ask me why if you're really curious ;) )

OpenStudy (anonymous):

okay yeah i actually knew that i swear hahaha. im confused cause since ihave to go down 3 AND up 3?? so would i have two different things ? like.... y-3=x^2 and y+3=x^2

terenzreignz (terenzreignz):

It's not down and up, it's down and right, LOL if it's going right, then you change the x-part, right?